OpenStudy (anonymous):

Beams for holding a sprinkler system are to be placed across the top of the greenhouse. The formula for the height h at which the beams are to be placed is given by.... Beams for holding a sprinkler system are to be placed across the top of the greenhouse. The formula for the height h at which the beams are to be placed is given by.... @Mathematics

OpenStudy (anonymous):

\[h=\sqrt{15^{2-}(a/2)^{2}}\]

OpenStudy (anonymous):

hey dint u figured it out??

OpenStudy (anonymous):

No i got stuck after i got 225-(a/2?^2)

OpenStudy (anonymous):

h^2 = 225 - a^2/4 => h^2 = (900 - a^2)/4 = [30^2 - a^2]/4 => h^2 = (30-a)(30+a)/4 => h = [(30^2-a^2)^(1/2)]/2

OpenStudy (anonymous):

Where did you get 900 from?

OpenStudy (anonymous):

\[h = \frac{\sqrt{30^2 - a^2}}{2}\]

OpenStudy (anonymous):

You set the problem up wrong I think. It looks like this on the paper... \[h=\sqrt{r ^{2}-(a/2)^{2}}\] ...... and r=15

OpenStudy (anonymous):

\[h = \sqrt{225 - \frac{a^2}{4}} = \sqrt{\frac{225 \times 4 - a^2}{4}} = \sqrt{\frac{900-a^2}{4}} = \frac{\sqrt{30^2 - a^2}}{2}\]

OpenStudy (anonymous):

\[r^2 = 15^2 = 15 \times 15 = 225\]

OpenStudy (anonymous):

now see these and say whether i set the wrong formulae??

OpenStudy (anonymous):

Ok so where do i go from there to solve for h

OpenStudy (anonymous):

if u will know the value of a then only u can solve it for h!!!

OpenStudy (anonymous):

a=25